OpenStudy (anonymous):

.Solve each system by graphing. 4. y = -4x y = - x + 6 (0, 6) (-6, 6) (-2, 8) (-8, 0)

hero (hero):

y = y -4x = -x+6 x - 6 = 4x -6 = 3x -6/3 = x -2 = x y = 8 (x,y) = (-2,8)

OpenStudy (anonymous):

You have to plot on a graph. choose two x ordinances for each equation. I choose x= 0, x= 1, you'll get the respective y's , (x, y) y=-4x and y= -x+ 6 ... (0,y1), (1, y2) you'll need two pts. for each line. Make sure you extend the line where they meet is the answer and it will correspond with Hero wrote, but you have to GRAPH as per the instructions.

OpenStudy (anonymous):

thanks